Calculate the temperature at which root mean square speed of gas molecules is double of its speed at 27°c pressure remaining constant

1
Expert's answer
2022-02-24T09:05:15-0500

"v=\sqrt{\frac{3RT}M},"

"v\sim \sqrt T,"

"T_2=2^2T_1=4T_1=1200~K=927\u00b0C."